OpenStudy (anonymous):

HELP PLEASE :( Calculate the moles of each element in C12H22O11. Calculate the number of atoms of each type in C12H22O11.

OpenStudy (aaronq):

to find moles, use the formula: \(\sf \huge moles=\dfrac{mass}{Molar~mass}\) to find atoms, use: \(\sf \huge moles=\dfrac{atoms}{Avogadro's~number}\) ----------------------- the molar mass of the compound is the sum of the individual molar masses of all the atoms in that compound. For \(C_{12}H_{22}O_{11}\) molar mass =12*Molar mass of C+22*molar mass of H+11*molar mass of O =12*12 g/mol+22*1 g/mol+11*16 g/mol ------------------------------